Pachyderm Trunk Co. has earnings and dividends that are expected to grow at an annual rate of 10% over the next 4 years. after that point earnings and dividends will grow consistently at 4% per year. current dividends are $1.40 per share. compute the value of Pachyderm stock for investors who require a 15% rate of return?
Answer #1

D1=(1.4*1.1)=1.54

D2=(1.54*1.1)=1.694

D3=(1.694*1.1)=1.8634

D4=(1.8634*1.1)=2.04974

Value after year 4=(D4*Growth rate)/(Required rate-Growth rate)

=(2.04974*1.04)/(0.15-0.04)

=19.37936

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.54/1.15+1.694/1.15^2+1.8634/1.15^3+2.04974/1.15^4+19.37936/1.15^4

which is equal to

=$16.10(Approx).